# VLog: Video-Language Models by Generative Retrieval of Narration Vocabulary

### v1

@@ Abstract
Human daily activities can be concisely narrated as sequences of routine events (e.g., turning off an alarm) in video streams, forming an event vocabulary. Motivated by this, we introduce VLog, a novel video understanding framework that define video narrations as vocabulary, going beyond the typical subword vocabularies in existing generative video-language models. Built on the lightweight language model GPT-2, VLog feature three key innovations: (i) A generative retrieval model, marrying language model’s complex reasoning capabilities with contrastive retrieval’s efficient similarity search. (ii) A hierarchical vocabulary derived from large-scale video narrations using our narration pair encoding algorithm, enabling efficient indexing of specific events (e.g., cutting a tomato) by identifying broader scenarios (e.g., kitchen) with expressive postfixes (e.g., by the left hand). (iii) A vocabulary update strategy leveraging generative models to extend the vocabulary for novel events encountered during inference. To validate our approach, we introduce VidCap-Eval, a development set requiring concise narrations with reasoning relationships (e.g., before and after). Experiments on EgoSchema, COIN, and HiREST further demonstrate the effectiveness of VLog, highlighting its ability to generate concise, contextually accurate, and efficient narrations, offering a novel perspective on video understanding. Codes are released at https://github.com/showlab/VLog.
@@ 1 Introduction
To leverage this narration vocabulary, a straightforward approach is to use a retrieval-based model for efficient narration retrieval. However, retrieval models often lack the reasoning capabilities of generative models, failing to handle advanced queries like “What’s the next action?” with a reference video. To address this issue, VLog introduces a novel generative retrieval architecture that repurposes the language model’s reasoning capabilities with a retrieval token, embedding both visual and query information for reasoning-oriented retrieval.
@@
  • We propose the first generative retrieval framework for video understanding, enabling efficient vocabulary indexing with support for complex reasoning.
  • We introduce the Narration Pair Encoding method and a hierarchical strategy to build the narration vocabulary and enable efficient indexing. To handle novel narrations beyond the existing vocabulary, we develop an efficient vocabulary expansion strategy.
@@ 2.2 Video-Language Models
Retrieval Models. Early vision-language models [39, 35, 4, 27, 50] primarily relied on alignment approaches, leveraging contrastive objectives for scalable performance. These methods excelled in retrieval and classification tasks within pre-defined labeled spaces, offering high efficiency by dot-product similarities calculation [25, 55, 38, 28]. But this fashion lacks the ability to model complex reasoning relationships between media inputs; for example, they fail to retrieve a video clip based on a query like, “What occurred after?” which conditioned on a reference video clip.
Generative Models. Recently, focus has shifted to generative LMMs, with several works [1, 23, 30, 61, 10, 24, 52] developing large multimodal models by projecting visual inputs into textual embeddings and aligning them with LLMs through visual instruction tuning. However, such straightforward adaptations are often unsuitable for video understanding, as large subword vocabularies in LLMs, while broadly inclusive, lack visual interpretability. Moreover, slow decoding during testing hinders real-time applications. Therefore, recent studies have focused on improving the efficiency of VideoLLM for (long) video modeling [44, 26, 8, 51, 57, 43].
Generative Retrieval. By examining the strengths and limitations of retrieval and generative models, a natural approach is to unify them, leveraging reasoning alongside efficient retrieval. In the realm of LLMs, efforts [48, 2, 45, 20] have been made to improve text embeddings that capture complex relationships, showing promise for practical applications. However, these advancements are less explored in multimodal domains [19, 31, 17], particularly in video understanding.
In this work, we pursue this generative retrieval direction by introducing a novel method that marrying lightweight GPT-2’s reasoning [40] with the contrastive vision-text model SigLIP [54]. Moreover, we departing from traditional subword vocabularies, VLog redefines the narration vocabulary, incorporating retrieval-based similarity search for fast and interpretable reasoning.
Figure 2: Comparison between different Video-Language model architectures: (a) Generative Models: These model with complex reasoning but are slow, generating tokens one by one. (b) Retrieval Models: These enable fast vocabulary search but lack reasoning, useful only for simple alignment tasks. (c) Generative Retrieval (VLog): This approach combines fast vocabulary search with complex reasoning by using a retrieval token, merging the advantages of both methods.
@@ 3.1 Architecture
This autoregressive process enables the model to capture complex dependencies between the visual inputs and the query, as demonstrated in Fig.2(a). However, for dense narrations over a long video, this approach incurs high inference costs due to token-by-token decoding [56, 49, 13].
@@
This approach offers two benefits: first, it enables efficient similarity calculation via dot product (i.e., Eq. 3); second, it directly outputs a narration without requiring subword generation. However, despite leveraging the strengths of contrastive models, this method struggle to capture the complex relationships in queries $\mathcal{Q}$, posing challenges for retrieving with a queries e.g., ‘What happened next.’
c. Generative Retrieval (Ours). To leverage the reasoning modeling of generative models while harnessing the efficiency of retrieval methods, we propose a novel generative retrieval model. As shown in Fig.2(c), we introduce a retrieval token with embedding $\mathbf{t}$ to bridge the generative language model and retrieval model. The retrieval token is positioned as the last input in the language model sequence, allowing it to attend to both the front visual and query inputs. After passing through the language model, the output embedding $\tilde{\mathbf{t}}$ is assumed to encode both visual and query information, enabling retrieval while preserving reasoning ability:
@@ 3.2 Vocabulary Construction
To enable model for prefix and postfix retrieval, we first use the retrieval token to retrieve the prefix, concatenate it with the visual and query sequence, and then use the retrieval token to retrieve a postfix from the postfix vocabulary.

### v2

@@ Abstract
Human daily activities can be concisely narrated as sequences of routine events (e.g., turning off an alarm) in video streams, forming an event vocabulary. Motivated by this, we introduce VLog, a novel video understanding framework that define video narrations as vocabulary, going beyond the typical subword vocabularies in existing generative video-language models. Built on the lightweight language model GPT-2, VLog feature three key innovations: (i) A generative retrieval model, marrying language model’s complex reasoning capabilities with contrastive retrieval’s flexible upgrading over narration vocabulary. (ii) A hierarchical vocabulary derived from large-scale video narrations using our narration pair encoding algorithm, enabling efficient indexing of specific events (e.g., cutting a tomato) by identifying broader scenarios (e.g., kitchen) with expressive postfixes (e.g., by the left hand). (iii) A vocabulary update strategy leveraging generative models to extend the vocabulary for novel events encountered during inference. To validate our approach, we introduce VidCap-Eval, a development set requiring concise narrations with reasoning relationships (e.g., before and after). Experiments on EgoSchema, COIN, and HiREST further demonstrate the effectiveness of VLog, highlighting its ability to generate concise, contextually accurate, and efficient narrations, offering a novel perspective on video understanding. Codes are released at https://github.com/showlab/VLog.
🖂: Corresponding Author.
@@ 1 Introduction
To leverage the narration vocabulary in LLMs, it need to train the embedding layers to accommodate newly added vocabulary units. In contrast, CLIP [39]-based retrieval models can directly extract embeddings for novel vocabulary without additional training, enabling efficient similarity matching. However, retrieval models often lack the reasoning capabilities of generative models, failing to handle advanced queries like “What’s the next action?” with a reference video. To address this issue, VLog introduces a novel generative retrieval architecture that repurposes the language model’s reasoning capabilities with a retrieval token, embedding both visual and query information for reasoning-oriented retrieval.
@@
  • We introduce the first narration vocabulary for VideoLLMs, which is constructed by our Narration Pair Encoding method and a hierarchical strategy to enable efficient indexing, resulting xxx speedup;
  • We introduce a novel generative retrieval strategy, which supports flexible vocabulary upgrading without training embedding for new vocabulary.
@@ 2.2 Video-Language Models
Retrieval Models. Early vision-language models [39, 35, 4, 27, 50] primarily relied on alignment approaches, leveraging contrastive objectives for scalable performance. These methods excelled in retrieval and classification tasks within pre-defined labeled spaces, offering high efficiency by narration-level dot-product similarities calculation [25, 55, 38, 28]. But this fashion lacks the ability to model complex reasoning relationships between media inputs; for example, they fail to retrieve a video clip based on a query like, “What occurred after?” which conditioned on a reference video clip.
Generative Models. Recently, focus has shifted to generative LMMs, with several works [1, 23, 30, 61, 10, 24, 52] developing large multimodal models by projecting visual inputs into textual embeddings and aligning them with LLMs through visual instruction tuning. However, such straightforward adaptations are often unsuitable for video understanding, as large subword-level vocabularies in LLMs, while broadly inclusive, lack visual interpretability. Moreover, slow decoding during testing hinders real-time applications. Therefore, recent studies have focused on improving the efficiency of VideoLLM for (long) video modeling [44, 26, 8, 51, 57, 43].
Generative Retrieval. By examining the strengths and limitations of retrieval and generative models, a natural approach is to unify them, leveraging reasoning alongside efficient retrieval. In the realm of LLMs, efforts [48, 2, 45, 20] have been made to improve text embeddings that capture complex relationships, showing promise for practical applications. However, these advancements are less explored in multimodal [19, 31, 17], particularly in video understanding.
In this work, we pursue this generative retrieval direction by introducing a novel method that marrying lightweight GPT-2’s reasoning [40] with the contrastive vision-text model SigLIP [54]. Moreover, we departing from traditional subword vocabularies, VLog redefines the narration vocabulary, incorporating retrieval-based search for flexible vocabulary upgrading and interpretable reasoning.
Figure 2: Comparison between different Video-Language model architectures: (a) Generative Models: These model with complex reasoning but are slow, generating subword tokens one by one, Vocabulary upgrading requires retraining. (b) Retrieval Models: These enable fast narration vocabulary via similarity matching and support vocabulary upgrading without training, but lack reasoning, useful only for simple alignment tasks. (c) Generative Retrieval (VLog): This approach combines fast narration-vocabulary and flexible vocabulary upgrading with complex reasoning by using a retrieval token, merging the advantages of both methods.
@@ 3.1 Architecture
This autoregressive process enables the model to capture complex dependencies between the visual inputs and the query, as demonstrated in Fig.2(a). However, for dense narrations over a long video, this approach incurs high inference costs due to token-by-token decoding [56, 49, 13]. Despite the efficiency gains brought by adopting the narration-level vocabulary, adding new vocabulary units to the embedding layer of LLMs typically requires training, making the approach less flexible.
@@
This approach offers two benefits: first, it enables narration-level retrieval via dot product similarity calculation (i.e., Eq. 3); second, the vocabulary is independent of LLM embedding weights, making it practical for upgrades when needed. However, despite leveraging the strengths of contrastive models, this method struggle to capture the complex relationships in queries $\mathcal{Q}$, posing challenges for retrieving with a queries e.g., ‘What happened next.’
c. Generative Retrieval (Ours). To leverage the reasoning modeling of generative models while harnessing the efficiency of narration-vocabulary and the flexiblity of vocabulary upgrading, we propose a novel generative retrieval model. As shown in Fig.2(c), we introduce a retrieval token with embedding $\mathbf{t}$ to bridge the generative language model and retrieval model. The retrieval token is positioned as the last input in the language model sequence, allowing it to attend to both the front visual and query inputs. After passing through the language model, the output embedding $\tilde{\mathbf{t}}$ is assumed to encode both visual and query information, enabling retrieval while preserving reasoning ability:
@@ 3.2 Vocabulary Construction
To enable model for prefix and postfix retrieval, we first use the retrieval token to retrieve the prefix, concatenate it with the visual and query sequence, then use the retrieval token to retrieve a postfix from the postfix vocabulary.
@@ 5 Conclusion and Limitations
Acknowledgements.
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).